OpenStudy (anonymous):

The hypotenuse of a right triangle is 25 cm long. One leg of the same triangle is 7 cm long. What is the length of the other leg? Explain how you found your answer.

OpenStudy (anonymous):

Just use the Pythagorean theorem.

OpenStudy (amistre64):

the answer is found by manipulating the pythag thrm by algebraic properties: 25^2 = 7^2 + s^2 25^2 - 7^2 = s^2 sqrt(25^2 -7^2) = s